OpenStudy (anonymous):

What is the approximate volume of the sphere? (Use 227 for π.) cubic inches

OpenStudy (jdoe0001):

well, is very obvious, since the sphere is "blank", its volume is "blank"

OpenStudy (triciaal):

volume of sphere = 4/3 * pi* r^3 use pi = 22/7 as soon as you figure out what is the radius, r, in inches, you can plug all the numbers in the formula to calculate the volume. unit of measure is cubic inches.